OpenStudy (anonymous):

five times a number k minus 9 is two thirds of a number j

OpenStudy (anonymous):

5k-9 = (2/3)j Now, we solve for j. Start, by multiplying both sides by 3: 15k-27 = 2j Now, divide both sides by 2: (15/2)k-(27/2) = j